A liquid take 10 minutes to cool from 90•c to 60•c. How much time will it take cool from 70•c to 30•c. The temperature of sorrounding is 30•c & Newton`s laws of cooling is applicable here

Grade:11

1 Answers

Adarsh
6 years ago
Dear Bhavesh,
according to newton law of cooling ,   T2-T1/t=K. {T1+T2/2 – TS}
Where T2 T1 are initial and final temp. t is time Ts  is surrounding temp. and k is any constant.
putting the value for cooling from 900C to 600C in 10 minute ,
90-60/10=k.{90+60/2 – 30} 
=> k=1/15 , now use this value of k for cooling from 700 to 30
70-30/t=1/15{50-30}
=> t = 30 minute
 
REGARDS 
Adarsh